# Intrinsic versus extrinsic cellular decision making

### v1

@@ Abstract
A cell routinely responds to one of many competing environmental cues. Does the cell have an intrinsic preference for that cue, or does that cue have the highest extrinsic information content? We introduce a theoretical framework to answer this fundamental question. We derive extrinsic detection limits for four types of directional cues—external and self-generated chemical gradients, fluid flow, and contact inhibition of locomotion—and thus predict extrinsic decision boundaries when these cues compete as pairs. Comparing the boundaries to published data from cell migration experiments quantitatively determines the degree to which cell decisions are intrinsic vs. extrinsic, revealing the extent of cells’ autonomy and providing interpretation of their response networks.
@@
Cells function in the presence of many competing environmental signals. These signals include chemical cues, mechanical cues, and cues from other cells. The traditional approach to understanding a cell’s response to an environmental cue focuses on characterizing the cell’s internal signaling network, which is ultimately responsible for transforming the cue into an observable behavior Hartwell et al. (1999); Alon (2019); Saha et al. (2022). A markedly different approach focuses on the information contained in the cue itself, investigating whether a cell’s response approaches that of an optimal physical detector Berg and Purcell (1977); Endres and Wingreen (2008); Bouffanais et al. (2013); Mugler et al. (2016); Fancher et al. (2020); Wang and Camley (2024). It is not clear which approach to use to understand a cell’s response to two or more competing cues: does the cell respond to the cue best matched to its intrinsic signaling network, or does it respond to the cue with the most extrinsic information? Here we make this fundamental question quantitative.
We focus on the ubiquitous response of cell migration. Migration is a cornerstone of tissue development Ewald et al. (2008), wound healing Phillipson and Kubes (2011), and disease progression Fenderson (2005), notably in cancer Friedl and Wolf (2003). Migration is guided by directional cues. Metastatic cancer cells, for instance, respond to chemical gradients Kim et al. (2013); Moon et al. (2021); ran Moon et al. (2023), including self-generated gradients Shields et al. (2007), as well as other directional cues such as pressure gradients generated by fluid flow Polacheck et al. (2011, 2014). Migratory cells also respond to each other’s presence through a process known as contact inhibition of locomotion, where cell-cell contact results in migration reversal Mayor and Carmona-Fontaine (2010); Lin et al. (2015). These cues often occur simultaneously, such as in complex tissue environments or tumors. This makes the question of responding to competing directional cues both fundamentally interesting and physiologically relevant.
@@
Here we investigate three experimentally motivated scenarios in which cells must respond to one of two competing cues (Fig. 1): (a) an autologous (self-generated) chemical gradient vs. a flow-induced pressure gradient, (b) an exogenous chemical gradient vs. a flow-induced pressure gradient, and (c) an exogenous chemical gradient vs. contact inhibition of locomotion from another cell. We first focus on the extrinsic information contained in each signal, deriving physical bounds on the precision with which each could be sensed. This approach leads to extrinsic decision boundaries separating phases where the cell is predicted to follow one cue over the other. Because each scenario corresponds to published experiments in which all free parameters have been measured, our theory makes falsifiable predictions that the published data either do or do not obey. Interestingly, we find that the extrinsic decision boundary is obeyed in case (a), partially obeyed in case (b), and violated in case (c). Therefore, we end by focusing on the implications of these findings for the intrinsic signaling network in each case. Throughout, our theory is built on simple scaling arguments because, when decision boundaries are violated, they are violated by orders of magnitude. This allows us to elucidate the rich physics underlying each sensing mode.
@@
Figure 1: The three experimentally motivated examples we investigate, where competing cues promote opposite migratory outcomes for a cell.
@@
Autologous chemical sensing vs. mechanical flow sensing. In the presence of a slow background fluid flow, multiple cancer cell types have been shown to secrete and detect a chemical called CCL21 whose distribution is biased by the flow Shields et al. (2007); Munson et al. (2013). This creates a downstream gradient that cells track, a process termed autologous chemotaxis Shields et al. (2007) . However, at high cell density, these cells instead migrate upstream due to mechanical sensing of the flow-induced pressure gradient across the cell body Polacheck et al. (2011). The hypothesis is that high cell density corresponds to a large background concentration of the secreted chemical, saturating the chemical gradient and allowing pressure sensing to take over Polacheck et al. (2011); Vennettilli et al. (2022). Thus far, this hypothesis has not been quantitatively tested.
@@
We begin by deriving the physical limit to the precision of chemical gradient sensing for this process. Throughout this work we take precision to be the signal-to-noise ratio for a physical quantity that contains the directional information of the cue. For a chemical gradient, it suffices to consider the difference $\Delta n=n_{2}-n_{2}$ in the number of molecules detected by the two halves of the cell [Fig. 2(a)] Mugler et al. (2016); Vennettilli et al. (2022); González and Mugler (2023). For slow flow, the mean of this difference scales as $\Delta\bar{n}\sim\epsilon\nu a^{2}/D$ Vennettilli et al. (2022), where $a$ is the cell lengthscale, $\nu$ is the molecule secretion rate, $D$ is the molecular diffusion coefficient, and the flow speed $v$ enters through the Peclét number $\epsilon=va/D\ll 1$ (see Table I in the Supplemental Material sup for a list of all parameters in this work, as well as their experimentally measured values). This expression for $\Delta\bar{n}$ follows from taking, in each half of the cell, the ratio of the rates of molecule gain due to secretion, and of loss due to diffusion and flow Vennettilli et al. (2022).
@@
Figure 2: Schematics of (a) autologous chemical sensing, (b) mechanical flow sensing, (c) exogenous chemical sensing, and (d) contact inhibition of locomotion. See text and Table I sup for parameter descriptions.
The variance in $\Delta n$ scales the same way as the variance in the total number $n$ of molecules detected by the cell Mugler et al. (2016). Because diffusion is Poissonian, this variance equals its mean $\bar{n}$. In a time $\tau$, the variance can be reduced as $\sigma^{2}_{\Delta n}\sim\bar{n}/M$ by the cell integrating $M=D\tau/a^{2}$ independent measurements, where $M$ follows from dividing $\tau$ by the diffusive correlation time $a^{2}/D$ Berg and Purcell (1977). The mean scales as $\bar{n}\sim\nu a^{2}/D+ca^{3}$, where the first term corresponds to molecules secreted by the cell, and the second term corresponds to a background concentration $c$ of molecules secreted by other cells Vennettilli et al. (2022). The background concentration is related to the cell density $\rho$ by balancing molecule secretion with molecule loss from flow, giving $c=(\nu L/v)\rho$, where $L$ is the length of the experimental environment in the flow direction Vennettilli et al. (2022).
This expression demonstrates that precision increases with faster flow, via $\epsilon$, but decreases with higher cell density $\rho$. Although it is derived using simple scaling arguments, these scalings are validated by the fact that expressions for $\Delta\bar{n}/\sigma_{\Delta n}$ with $\rho=0$, and for $\Delta\bar{n}/\bar{n}$ with $\rho>0$, agree  [^1] Specifically, $\Delta\bar{n}/\sigma_{\Delta n}=\epsilon\sqrt{\nu\tau}$ with $\rho=0$ agrees with Eq. 15 in Fancher et al. (2020), and $\Delta\bar{n}/\bar{n}=\epsilon^{2}/(\epsilon+\rho a^{2}L)$ agrees with Eq. 8 in Vennettilli et al. (2022), up to factors of order unity. with more rigorous results derived by us previously using a reaction diffusion treatment that accounts explicitly for the fluid flow field Fancher et al. (2020); Vennettilli et al. (2022).
We now seek to compare this result with the physical limit to the precision of mechanical sensing of the flow. Previous work derived this limit for cells detecting a flow via mechanosensitive ion channels that respond to changes in membrane tension when the flow induces a shear Bouffanais et al. (2013). However, when cells are embedded in a dense extracellular matrix, as in the experiments Polacheck et al. (2011, 2014), the flow is low-permeability, meaning that the average pore size of the matrix is much smaller than the cell. At low permeability, the force on the cell due to the flow is dominated by pressure, not shear Polacheck et al. (2014). Therefore we adapt the previous treatment to pressure sensing. Again, we focus on simple scaling arguments, which to our knowledge are absent from the literature for this process.
The flow induces a pressure difference $\Delta P$ across the cell. At low permeability $K\ll a^{2}$, the pressure difference scales with the flow velocity $v$ as $\Delta P\sim v\eta a/K$ Polacheck et al. (2014), where $\eta$ is the fluid viscosity. The pressure difference causes a membrane tension difference via the Young-Laplace equation, $\Delta\gamma\sim a\Delta P$. As in the previous work Bouffanais et al. (2013), we suppose that the cell detects this tension difference using $N$ ion channels distributed across the membrane [Fig. 2(b)]. Each channel switches between an open and a closed conformation. The probability $p$ that the channel is open obeys that for a two-state system in a thermal bath, $p=(1+e^{-q})^{-1}$, where $q=\beta\Delta E$ is the energy difference $\Delta E$ between the closed and open states, scaled by the inverse temperature $\beta$. The energy difference is given by $\Delta E=\Delta\gamma\Delta A$, where $\Delta A$ is the difference in channel area between the closed and open state Bouffanais et al. (2013); Ursell et al. (2008). Thus, an increase in pressure on the upstream side of the cell causes an increase in membrane tension, which increases the probability that a channel there is open.
Figure 3: Predicted extrinsic decision boundaries (color maps) and observed cell behavior (data points) for the three cases in Fig. 1. Data from Polacheck et al. (2011) (a, circles), Polacheck et al. (2014) (a, square), ran Moon et al. (2023) (b), and Lin et al. (2015) (c). See sup for analysis details and cod for code.
Just as the directional information of the chemical cue was contained in the difference in the detected molecule numbers from the two halves of the cell, here the directional information of the mechanical cue is contained in the difference in the (dimensionless) tension readouts $q$ from channels on the two halves of the cell. On the upstream half the $N/2$ channels will read out a tension increase, whereas on the downstream half the $N/2$ channels will read out a tension decrease. Thus, the mean difference will scale as $\bar{Q}\sim(N/2)q-(N/2)(-q)=Nq$. The variance, assuming the channels are sufficiently independent Bouffanais et al. (2013), will scale as $\sigma_{Q}^{2}\sim N\sigma_{q}^{2}/M$, where here $M=(k_{+}+k_{-})\tau$ because the correlation time is the inverse sum of the channel switching rates $k_{+}$ and $k_{-}$. The variance in $q$ propagates in quadrature from $p=(1+e^{-q})^{-1}$ as $\sigma_{q}^{2}=1/\sigma_{p}^{2}$, where $\sigma_{p}^{2}=\bar{p}(1-\bar{p})$ for a two-state switch. The minimum noise occurs when the channels have $\bar{p}=1/2$ (and thus $k_{+}=k_{-}\equiv k$) in the absence of flow, a condition called critical pre-stress Bouffanais et al. (2013).
This expression demonstrates that precision increases with faster flow, via $q$. Again, although it is derived using simple scaling arguments, it agrees with the previous work Bouffanais et al. (2013) in the case that tension change is induced by shear instead of pressure  [^2] Specifically, Eq. 2 agrees with Eqs. 41 and 42 in Bouffanais et al. (2013) up to factors of order unity when (i) the tension change $\Delta\gamma=a\eta G$ arises from the shear rate $G$ rather than the pressure difference, (ii) critical pre-stress is assumed ($\Delta h-\gamma_{0}\Delta A=0$ there), and (iii) the membrane stretch energy (much smaller than the tension energy) is ignored. Note that $k$ here then corresponds to $k_{0}e^{-\beta\gamma_{0}\Delta A/2}$ there, which given the parameter choices there evaluates to about once per second, consistent with the measured value we use here (Table I sup )..
All parameters in Eqs. 1 and 2 have been measured experimentally (Table I sup ). Therefore, we can evaluate the ratio ${\cal P}_{\rm chem}^{\rm auto}/{\cal P}_{\rm mech}$ with no free parameters. Figure 3(a) shows this ratio as a function of flow speed and cell density (color map). We see that, as hypothesized, the precision of mechanical sensing overtakes that of chemical sensing as the cell density increases. Overlaid as data points are the measured flow speeds and cell densities of cells that are observed to travel upstream or downstream Polacheck et al. (2011, 2014) (see sup for details on how all data points in Fig. 3 are extracted). The data approximately agree with the decision boundary predicted by the precision ratio. This finding suggests that these cells decide between autologous chemical sensing and mechanical flow sensing based on extrinsic signal information.
Exogenous chemical sensing vs. mechanical flow sensing. We previously investigated a cancer cell line exposed to both a fluid flow and a gradient of an exogenous chemical attractant TGF-$\beta$ ran Moon et al. (2023). We found that in locations where the gradient was downstream and steep, cells migrated downstream. However, in locations where the gradient was downstream but shallow, cells migrated upstream. Here we ask whether these observations are consistent with the precision crossover between exogenous chemical sensing and mechanical flow sensing.
Our derivation for exogenous chemical sensing follows that above for autologous chemical sensing, but now the average number $\bar{n}$ of detected molecules and the difference $\Delta\bar{n}$ between the two halves of the cell are set by the local background concentration $c$ and gradient $g$ of an exogenous chemical signal [Fig. 2(c)]. Specifically, we have $\bar{n}\sim ca^{3}$ and $\Delta\bar{n}\sim(ga)a^{3}$, where $ga$ is the approximate concentration difference across the cell Mugler et al. (2016). The other elements follow from above: $\sigma^{2}_{\Delta n}\sim\bar{n}/M$, with $M=D\tau/a^{2}$. Thus we arrive at the precision of exogenous chemical sensing,
The dimensionless parameter $G$ captures the fact that precision increases with the gradient but decreases with the background concentration: it is harder to measure a molecule number difference against larger background fluctuations Mugler et al. (2016). Once again, although this expression is derived using simple scaling arguments, it agrees  [^3] Specifically, Eq. 3 agrees with Eq. 11 in Mugler et al. (2016) up to factors of order unity in the limit of a single cell ($n_{0}=1$ there). with a more rigorous result derived by us previously using a reaction diffusion treatment that accounts for internal signaling within the detector Mugler et al. (2016).
Again, because all parameters in Eqs. 2 and 3 have been measured experimentally (Table I sup ), we can evaluate the ratio ${\cal P}_{\rm chem}^{\rm exo}/{\cal P}_{\rm mech}$ with no free parameters. Figure 3(b) shows this ratio as a function of flow speed and dimensionless gradient strength $G$ (color map). We see that chemical sensing is more precise for larger gradients, while mechanical sensing is more precise for larger flow speeds. Overlaid as data points are the measured flow speeds and $G$ values from our previous experiments ran Moon et al. (2023). Most points agree with the decision boundary predicted by the precision ratio, with one point pushing the boundary but still within error. This suggests that in most cases the decision is based on extrinsic information. However, the notable exception is the white point deep within the blue regime: these cells are predicted to migrate upstream but are observed, on average, to migrate nowhere.
The difference between the white data point and the other data points in Fig. 3(b) is that the white point corresponds to a very high background concentration (see sup ). In fact, here the gradient even points slightly upstream ($g<0$), meaning that, for this point only, both cues promote upstream migration with no conflict. Yet, the high background concentration evidently prevents cells from responding to either cue. This is in marked contrast to the previous case [Fig. 3(a)], where, like here, the high background concentration reduced the precision of chemical sensing below that of mechanical sensing, but unlike here, the cells indeed migrated upstream. This contrast has important implications for the intrinsic signaling networks in these two cases, which we expand upon in the Discussion.
Exogenous chemical sensing vs. contact inhibition of locomotion. Previous work investigated a cancer cell line exposed to a gradient of an exogenous chemical attractant EGF and placed in narrow channels Lin et al. (2015). This setup caused many encounters where up-gradient migration conflicted with repulsion from another cell due to contact inhibition of locomotion (CIL). At low background EGF concentration, cells migrated up-gradient, whereas at high background concentration, cells migrated away from the other cell. Here we ask whether these observations are consistent with the precision crossover between exogenous chemical sensing and CIL.
Recent work investigated the physical aspects of CIL accuracy Wang and Camley (2024) but focused on the effects of contact size and signal interference rather than a general precision limit. Here, as above, we derive the physical limit from simple scaling arguments. The mechanism of CIL is that each cell’s membrane has molecules called ephrin, as well as receptors to detect that molecule on another cell Lin et al. (2015); Wang and Camley (2024) [Fig. 2(d)]. Calling the contact area $A_{2}$ and the two-dimensional concentration of ephrin $c_{2}$, the mean number of ephrin moelcules scales as $\bar{n}\sim c_{2}A_{2}$. Calling the diffusion coefficient of ephrin $D_{2}$, the variance in this number, as above, scales as $\sigma_{n}^{2}\sim\bar{n}/M$ with $M=D_{2}\tau/A_{2}$. Thus we arrive at the precision of CIL,
${\cal P}_{\rm CIL}=\frac{\bar{n}}{\sigma_{n}}\sim\sqrt{D_{2}c_{2}\tau}.$ (4)
We note that the contact area $A_{2}$ drops out, which is consistent with the finding in the previous work that accuracy is largely insensitive to contact size Wang and Camley (2024).
Figure 3(c) shows the ratio ${\cal P}_{\rm chem}^{\rm exo}/{\cal P}_{\rm CIL}$ as a function of ephrin turnover $D_{2}c_{2}$ and dimensionless gradient strength $G$ (color map), again with all remaining parameters measured experimentally (Table I sup ). We see that chemical sensing is more precise for larger gradients, while CIL is more precise for higher ephrin turnover. Overlaid as data points are the measured $G$ values from the narrow channel experiments Lin et al. (2015) and the values of $D_{2}c_{2}$ from experiments that measured the diffusion and concentration of ephrin Xu et al. (2011). We see that the measured decision boundary is three orders of magnitude away from the predicted boundary. This finding suggests that these cells do not decide between exogenous chemical sensing and CIL based on extrinsic signal information alone. Specifically, it suggests that cells intrinsically weight information about the EGF gradient over information from CIL.
Discussion. What are the implications of Fig. 3 for the intrinsic signaling network in each case? The strong violation of the extrinsic decision boundary in Fig. 3(c) suggests that the EGF pathway and the CIL pathway interact upstream of the migration response, rather than transmitting the extrinsic information to the migration response independently. In fact, in the previous experimental work on this case Lin et al. (2015), it was demonstrated that the ephrin receptor represses the activity of the EGF receptor directly, far upstream of the cascade that connects this activity to the migration response [Fig. 4(c)]. The cell could easily tune the strength of this repression, with weaker repression suppressing information from the CIL signal and shifting the decision boundary further into the naively CIL-dominated regime, as in Fig. 3(c).
Figure 4: Pathway structure (a) that we predict or (b, c) that is previously known from experiments. (d) The three cases define a spectrum from extrinsic to intrinsic decision-making.
Similarly, the violation of the extrinsic decision boundary at high background concentration in Fig. 3(b) (white point) but not Fig. 3(a) suggests that the TGF-$\beta$ pathway and the flow sensing pathway converge upstream of the migration response. In fact, in our previous experimental work, we found that the results in Fig. 3(b), as well as many other results including cases of commensurate (rather than conflicting) cues, are explained by a model in which the two pathways converge on a common network component whose activity can saturate Moon et al. (2021); Saha et al. (2022); ran Moon et al. (2023) [Fig. 4(b)]. Conversely, while to our knowledge the interaction between the CCL21 pathway and the flow sensing pathway is less characterized, our results in Fig. 3(a) predict that these two pathways transmit the extrinsic information to the migration response independently [Fig. 4(a)]. Taken together, our results imply a spectrum of extrinsic to intrinsic decision-making for these three cases [Fig. 4(d)], with cases (a), (b), and (c) progressively more intrinsic, respectively.
By focusing on the physics of sensory precision, we have made quantitative and falsifiable the question of how cells respond to competing cues. Our approach can be extended straightforwardly to other cue types, to alternative cue pairings, and to competitions among more than two cues, or more than two responses. Our framework predicts the signals that cells transmit faithfully, and the signals that cells may suppress or ignore in favor of others. It facilitates the interpretation of signaling networks and may guide the discovery of pathways that are still poorly characterized.
Acknowledgements.
We thank Hye-ran Moon and Andre Levchenko for helpful discussions. The code for Fig. 3 is freely available cod . References Fleury et al. (2006); Sukharev et al. (1999); Morris (1990); Shapovalov and Lester (2004) are in the Supplemental Material sup . L. G. and A. M. were supported by National Science Foundation Grants No. PHY-2118561 and No. MCB-2118037. B. H. was supported by National Institutes of Health Grants No. R01 CA254110, No. R33 HL159948, and No. U01 CA274304.
SUPPLEMENTAL MATERIAL
I Parameter values for theory
$a$ Cell lengthscale 10 $\mu$m Polacheck et al. (2011, 2014); ran Moon et al. (2023); Lin et al. (2015)
$\nu$ Molecule secretion 1 s-1 Shields et al. (2007)
$D$ Molecule diffusion 150 $\mu$m2/s Fleury et al. (2006)
$L$ Chamber length 3 mm Polacheck et al. (2011)
$K$ Permeability $0.1$ $\mu$m2 Polacheck et al. (2011) Fig. 3(a)
$0.05$ $\mu$m2 ran Moon et al. (2023) Fig. 3(b)
$\Delta A$ Channel area change $6.5$ nm2 Sukharev et al. (1999)
$N$ Channels per cell $1000$ Morris (1990)
$k$ Channel switching rate $1$ s-1 Shapovalov and Lester (2004)
$\rho$ Cell density Fig. 3(a) Polacheck et al. (2011, 2014)
$v$ Fluid flow speed Fig. 3(a, b) Polacheck et al. (2011, 2014); ran Moon et al. (2023)
$c$ Background conc. Fig. 3(b, c) ran Moon et al. (2023); Lin et al. (2015)
$g$ Concentration gradient Fig. 3(b, c) ran Moon et al. (2023); Lin et al. (2015)
$c_{2}$ Ephrin concentration Fig. 3(c) Xu et al. (2011)
$D_{2}$ Ephrin diffusion Fig. 3(c) Xu et al. (2011)
@@ I Parameter values for theory
$\beta$ Inverse temperature $2.5\times 10^{20}$ J-1 Room temp.

### v2

# Quantifying cellular autonomy in multi-cue environments

@@ Abstract
A cell routinely responds to one of many competing environmental cues. A fundamental question is whether the cell follows the cue prioritized by its internal signaling network or the cue that carries the most external information. We introduce a theoretical framework to answer this question. We derive information limits for four types of directional cues: external and self-generated chemical gradients, fluid flow, and contact inhibition of locomotion. When the cues compete as pairs, these limits predict which cue a cell should follow if its decision is based on environmental information alone. We compare these predicted decision boundaries with data from our and others’ cell migration experiments, finding cases where the boundary is obeyed and cases where it is violated by orders of magnitude. Both outcomes are informative, and we find that they rationalize known properties, or predict putative properties, of cells’ internal signaling networks. Our work introduces a physical framework to quantify the degree to which a cell acts like an autonomous agent, rather than a passive detector, favoring a cue even when it is less informative.
@@
I Introduction
Cells navigate their environment by interpreting multiple, often competing, signals—chemical gradients, mechanical forces, and interactions with other cells. A key question is how cells integrate these cues to make decisions. One approach emphasizes the cell’s signaling network, which determines how a cue is processed internally [1, 2, 3]. Another approach considers the external information available in each cue, asking whether a cell’s behavior approximates that of an optimal physical detector [4, 5, 6, 7, 8, 9]. When signals conflict, does a cell follow the cue prioritized by its signaling network, or the cue that carries the most information? Here, we make this question quantitative and test it experimentally.
We focus on directed cell migration, a fundamental process in development [10], wound healing [11], and cancer metastasis [12]. Migrating cells respond to chemical gradients [13, 14, 15], including self-generated ones [16], as well as mechanical signals like shear [17, 18] or pressure gradients [19, 20] induced by fluid flow. They also interact with other cells through contact inhibition of locomotion, where contact reverses migration [21, 22]. In complex environments such as tumors, multiple cues are present simultaneously, making the problem of cue selection both fundamental and physiologically relevant.
@@ I Introduction
We investigate three experimental scenarios in which cells must respond to one of two competing cues (Fig. 1): (a) an autologous (self-generated) chemical gradient vs. a flow-induced pressure gradient, (b) an exogenous chemical gradient vs. a flow-induced pressure gradient, and (c) an exogenous chemical gradient vs. contact inhibition of locomotion from another cell. We consider the information contained in each signal, deriving physical bounds on the precision with which each could be sensed. This approach leads to decision boundaries separating phases where the cell is predicted to follow one cue over the other if the cell optimally detects and equally weighs the cues in its decision. We then compare these predictions with data from experiments, both our own and previously published. Because all free parameters of the theory have been measured in these experiments, the theory makes falsifiable predictions that the data either do or do not obey.
@@
Figure 1: The three experimental examples we investigate, where competing cues promote opposite migratory outcomes for a cell.
We find that the decision boundary is obeyed in case (a), partially obeyed in case (b), and strongly violated in case (c). These findings have important implications for the cellular decision-making process in each case. In case (a), for example, the cell’s decision is equivalent to that of a passive, comparative detector, whereas in case (c) the cell’s decision reflects a strong preference for a one cue over another, even when it has the smaller information content. We map these outcomes to the known features of the cell’s signaling network in each case. Where these features are known in detail, our findings provides rationales for their function. Where they are not known, our findings make predictions for their structure. Altogether, our results introduce and demonstrate a framework for quantifying cellular autonomy: the degree to which a cell favors a cue even when it is less informative.
II Results
We begin by defining the detectable information in four types of environmental cues and deriving physical limits to the detection precision. In each case, we take precision to be the signal-to-noise ratio for a physical quantity that contains the directional information of the cue. Our theory is built on simple scaling arguments because we find that when decision boundaries are violated, they are violated by orders of magnitude. In all cases, we will find that our results agree with previous results derived more rigorously but in special cases. This validates our scaling approach, even as we go beyond these cases. Throughout this work we assume spherical cells with uniform, noninteracting sensors, and we consider relaxations of these assumptions in the Discussion.
II.1 Autologous chemical sensing vs. mechanical flow sensing
@@ II Results
In the presence of a slow background fluid flow, multiple cancer cell types have been shown to secrete and detect a chemical called CCL21 whose distribution is biased by the flow [16, 23]. This creates a downstream gradient that cells track, a process termed autologous chemotaxis [16]. However, at high cell density, these cells instead migrate upstream due to mechanical sensing of the flow-induced pressure gradient across the cell body [19]. The hypothesis is that high cell density corresponds to a large background concentration of the secreted chemical, saturating the chemical gradient and allowing pressure sensing to take over [19, 24]. Despite important theoretical work on this problem [24, 25, 26], this hypothesis has not been quantitatively tested.
II.1.1 Autologous chemical sensing
@@
We first derive the physical limit to the precision of chemical gradient sensing for this process. The directional information is contained in the difference $\Delta n=n_{2}-n_{1}$ in the number of molecules that could be detected by the two halves of the cell [Fig. 2(a)] [7, 24, 27]. For slow flow, the mean of this difference scales as $\Delta\bar{n}\sim\epsilon\nu a^{2}/D$ [24], where $a$ is the cell lengthscale, $\nu$ is the molecule secretion rate, $D$ is the molecular diffusion coefficient, and the flow speed $v$ enters through the Peclét number $\epsilon=va/D\ll 1$ (see Table I for a list of all parameters in this work, as well as their experimentally measured values). This expression for $\Delta\bar{n}$ follows from taking, in each half of the cell, the ratio of the rate of molecule gain due to secretion, to the rate of loss due to diffusion and flow [24].
@@
Figure 2: Schematics of (a) autologous chemical sensing, (b) mechanical flow sensing, (c) exogenous chemical sensing, and (d) contact inhibition of locomotion. See text and Table I for parameter descriptions.
Meaning Value Reference
$a$ Cell lengthscale 10 $\mu$m [19, 20, 15, 22]
$\nu$ Molecule secretion 1 s-1 [16]: Fig 3(a) main
$6$ s-1 [23]: Fig 3(a) inset
$D$ Molecule diffusion 150 $\mu$m2/s [28, 23]
$L$ Chamber length 3 mm [19, 29]
$K$ Permeability $0.1$ $\mu$m2 [19]: Fig 3(a)
$0.05$ $\mu$m2 [15]: Fig 3(b)
$\Delta A$ Channel area change $6.5$ nm2 [30]
$N$ Channels per cell $1000$ [31]
$k$ Channel switching rate $1$ s-1 [32]
$\rho$ Cell density Fig 3(a) [19, 20, 29, 23]
$v$ Fluid flow speed Fig 3(a, b) [19, 20, 29, 23, 15],
this work
$c$ Background conc. Fig 3(b, c) [15, 22], this work
$g$ Concentration gradient Fig 3(b, c) [15, 22], this work
$c_{2}$ Ephrin concentration Fig 3(c) [33]
$D_{2}$ Ephrin diffusion Fig 3(c) [33]
@@
$\beta$ Inverse temperature $2.5$$\times$$10^{20}$/J Room temp.